Bodin Company manufactures finger splints for kids who get tendonitis from playing video games. The firm had the following inventories at the beginning and end of the month of January.

January 1 January 31
Finished goods $ 125,000 $ 117,000
Work in process 235,000 251,000
Raw material 134,000 124,000

The following additional data pertain to January operations.

Raw material purchased $ 191,000
Direct labor 350,000
Actual manufacturing overhead 170,000
Actual selling and administrative expenses 115,000


The company applies manufacturing overhead at the rate of 60 percent of direct-labor cost. Any overapplied or underapplied manufacturing overhead is accumulated until the end of the year.

Exercise 3-27 Part 1

Required:

1. Compute the company’s prime cost for January.

prime cost=

2. Compute the total manufacturing cost for January.

total manufacturing cost=

3. Compute the cost of goods manufactured for January.

cost of goods manufactured=

4. Compute the cost of goods sold for January.

cost of goods sold=

5. Compute the balance in the manufacturing overhead account on January 31. Debit or credit?

Manufacturing overhead account balance on January 31 is
Answer #1

1) Direct material used = 134000+191000-124000 = 201000

Prime cost = Direct material+Direct labor = 201000+350000 = 551000

2) Total manufacturing cost = 551000+350000*60% = 761000

3) Cost of goods manufactured = 235000+761000-251000 = 745000

4) Cost of goods sold = 125000+745000-117000 = 753000

5) Applied overhead = 350000*60% = 210000

Actual overhead = 170000

Manufacturing overhead balance = $40000 Credit